OpenStudy (anonymous):

A sample of four ball is to be selected at random from an urn containing 16 balls numbered 1 to 16. If seven of the balls are green, four are white, and five are black, how many different samples of four balls contain all three colours?

OpenStudy (anonymous):

7 x 4 x 5 x 13

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

the answer is 910. I don't know how to solve.

OpenStudy (anonymous):

ok, so my answer is pretty close, only off by a factor of 2 , so now I have to think why

OpenStudy (valpey):

gwbg = 7*4*5*6/2 +gwbw = 7*4*5*3/2 +gwbb = 7*4*5*4/2 =7*4*5*(6+3+4)/2

OpenStudy (valpey):

divide by 2 because the order of the two same colored balls doesn't matter